OpenStudy (anonymous):

write the equation of the line that contains (1,4) and (3,10) in point slope, slope intercept, and standard form

OpenStudy (jim766):

find slope y2 - y1 ------ = m x2-x1 point slope y - y1 = m(x - x1) you put in the y1, the x1 and the m

OpenStudy (jim766):

x1 y1 x2 y2 ( 1, 4) (3, 10)

OpenStudy (anonymous):

We can write : \[y=ax+b\] where : \[a=\frac{y_2-y_1}{x_2-x_1}=\frac{10-4}{3-1}=3\] For finding b, because (1,4) is from the line, we have : \[4=1\times a+b=3+b\] so : \[b=4-3=1\] Hence , the equation is : \[y=3x+1\]

OpenStudy (anonymous):

Whats the m, is it the slope?

OpenStudy (anonymous):

and for the point slope does it matter which set of variables you use

OpenStudy (jim766):

if you mean, does it matter which points you use? no you can use either one

OpenStudy (anonymous):

cool beans

OpenStudy (anonymous):

how do I turn this into standard form?

OpenStudy (jim766):

start with y = 3x +1 now get the 1 by itself

OpenStudy (jim766):

now the trick with standard from is that the x's go first and they have to be positive what did you do to get the 1 by itself?

OpenStudy (jim766):

or are you still working on the point slope?

OpenStudy (jim766):

you good here?